OpenStudy (anonymous):

2k2 + 6k + 180 = (k - 6)(k + __). To complete this equation is 12? Need help please

OpenStudy (anonymous):

\[2k^2 + 6k + 180 = (k - 6)(k + ?)\] Find the roots of \[2k^2 + 6k + 180 = 0 \]

OpenStudy (anonymous):

so still need help

OpenStudy (anonymous):

\[2k^2 + 6k + 180 = (k - 6)(k + a)\] \[2k^2 + 6k + 180 = (k^2 - 6k + ak + 6a)\] you can solve by this method

OpenStudy (triciaal):

method ok but that should be a -6a at the end not +6a

OpenStudy (anonymous):

that is what its showing that I posted :(

OpenStudy (triciaal):

I meant what was just above my post

OpenStudy (anonymous):

im coming up with false but wanna make sure that its not 12 to finish the equation

OpenStudy (triciaal):

(k-6)(k + a) the last entry when expanded (-6)*(+a) = - 6a

OpenStudy (anonymous):

ok thankyou so much :)

OpenStudy (anonymous):

I don't have a -6a:(

OpenStudy (anonymous):

36 -5 9 -9 these are the answers